### v3

@@ 3.4 Training Strategy
The UVZ training process is supervised in two stages. In the first stage, we compute losses for the outputs of DEN and ASN, respectively, which is an offline pre-training process. During the second stage, we use DEN with fixed parameters to assist DGEN in generating the enhanced image, thereby obtaining relevant losses. The first stage of training is conducted by synthesized dataset [41], which contains underwater images and corresponding depth maps of five open ocean waters and coastal waters. We randomly selected 600 image pairs to train DEN and ASN for 100 epochs, and calculated the following $l_{1}$ loss:
@@
In the second stage, we selected 1200 image pairs from the CYCLE [42] dataset, with 1080 pairs used for 100 epochs training and the remaining 120 pairs used for subsequent testing. To ensure the enhanced images with pleasing colors and details, we formulate the total loss $L_{s}$ of DGEN using charbonnier loss and ssim loss. The former measures pixel similarity, while the latter evaluates structural similarity. The formula is as follows:
@@ 4.1 Implementation Details
To assess the performance of UVZ, we use the remaining 120 image pairs from CYCLE dataset for comparison experiments. Additionally, we used four benchmark datasets: LNRUD [43], LSUI [44], UFO [45], and UIEB [9]. We randomly selected 500 images from LNRUD, 400 images from LSUI, 120 images from UFO, and 80 images from UIEB for more comprehensive testing.
The proposed model was trained and tested on NVIDIA GeForce RTX 2080 Ti GPUs using the PyTorch framework. We initialized the learning rate to 0.0002 and adjusted it gradually with Adam optimizer. For the CYCLE dataset, we measure the similarity between the enhanced results and the reference images by three common referenced metrics (i.e., PSNR, SSIM and MSE). For all datasets, we evaluate the performance by three no-reference metrics (i.e., UIQM [46], UICM and NIQE).
@@ 4.2 Comparison on reference dataset
Our UVZ was compared with some state-of-the-art methods, including four traditional methods (i.e., L2UWE [5], CBM [6], TEBCF [7], WWPE [8]), four deep learning methods (i.e., Water-Net [9], SCNet [10], TACL [11], Spectroformer [14]), and two depth-related methods (i.e., USUIR [12], PUGAN [13]).
Figure 6: Enhancement comparison on the CYCLE dataset. From left to right are the raw images, the results of L2UWE [5], CBM [6], TEBCF [7], WWPE [8], and Water-Net [9], SCNet [10], TACL [11], USUIR [12], PUGAN [13], Spectroformer [14], the proposed UVZ, and the ground truth. The red box represents the enlarged details.
@@
Fig. 6 shows the enhancement results of all methods on the CYCLE dataset. L2UWE and USUIR improve image contrast, but noticeable color deviations persist. CBM and Water-Net have corrected colors to some extent, but the enlarged image details are blurred. TEBCF significantly enhances the details of the image but results in darkened brightness and color confusion. WWPE, TACL, PUGAN, and Spectraformer alleviate the effects of color deviation, but enlarged details are still affected by blue-green. SCNet and our UVZ effectively eliminate color deviation and enhance image visibility, with the latter highlighting more vibrant colors. Moreover, we also exhibit the histograms of the red channel in Fig. 7, which is the most severely lost during underwater imaging. The curves of UVZ and GT overlap the most, indicating the closest similarity. To further observe the color distribution, we show the pixel statistics on the RGB channels in Fig. 8. All methods enhance the degraded red channel to some extent, with UVZ being the most prominent. Its distribution on the red, green, and blue channels is 36.3%, 36.1%, and 29.2% respectively, which is more consistent with GT’s 35.5%, 34.5%, and 29.9%.
@@
L2UWE CBM TEBCF WWPE Water-Net SCNet TACL USUIR PUGAN
Spectro-
former
UVZ
PSNR↑ 14.55 21.48 18.74 17.40 23.23 22.48 21.78 20.55 21.43 21.42 26.00
SSIM↑ 0.6440 0.7461 0.6839 0.6412 0.8196 0.8116 0.7701 0.7739 0.7790 0.7910 0.8721
MSE↓ 0.06195 0.01745 0.02090 0.02560 0.01150 0.01081 0.01597 0.02211 0.01571 0.01639 0.00435
UIQM↑ 2.634 4.151 4.606 4.291 4.573 4.629 4.444 4.321 4.522 4.332 5.044
UICM↑ -76.35 -31.83 -13.71 -16.63 -16.43 -11.01 -23.29 -21.15 -20.74 -20.74 4.781
NIQE↑ 5.045 6.548 5.137 6.387 5.310 5.203 4.325 4.941 5.253 5.017 6.586
Runtime(s) 3.6725 0.2425 1.6340 0.3207 0.9613 0.0520 0.0595 0.0255 0.0488 0.0394 0.0583
@@ 4.3 Comparison on no-reference dataset
We present enhancement examples from LNRUD, LSUI, UFO and UIEB datasets in Figs. 9-12. In Fig. 9 and Fig. 10, the color enhancement from the UVZ is the most pronounced and well-established. While other methods mitigate the effects of color deviation, blue-green color still dominates the results. We enlarged the fishes in the near-far scenes in Fig. 9 (c) and Fig. 10 (c), respectively. It can see that only TEBCF and UVZ effectively remove the blue-green color, but the dull color of the former reduces the attractiveness.
From Fig. 11 and Fig. 12, it can be observed that L2UWE exhibits bright contrast enhancement, but over-exposure affects the color. The enhanced results of TEBCF, WWPE and Water-Net have low brightness and poor clarity. CBM effectively enhances image details, while TACL, USUIR and PUGAN alleviate the impact of color deviation. However, blue-green colors still dominate in their enhanced images. Despite the significant saturation increase in SCNet and Spectroformer, they still show deviated colors in the far scenes. Benefiting from its depth perception ability, our UVZ has excellent performance in terms of color, saturation, and visibility, whether in near or far scenes.
@@
L2UWE CBM TEBCF WWPE Water-Net SCNet TACL USUIR PUGAN
Spectro-
former
UVZ
LNRUD UIQM↑ 2.688 4.037 4.436 4.296 4.442 4.452 4.326 4.296 4.377 4.314 4.619
UICM↑ -74.80 -31.29 -16.69 -17.15 -19.72 -16.85 -27.65 -21.25 -23.03 -20.33 -9.933
NIQE↑ 5.322 6.684 5.485 5.550 5.694 5.305 4.465 5.374 5.133 5.308 6.241
LSUI UIQM↑ 2.882 4.215 4.635 4.439 4.607 4.597 4.460 4.672 4.542 4.881 4.863
UICM↑ -72.28 -29.06 -14.98 -13.64 -15.31 -11.81 -22.39 -20.31 -17.48 -15.64 -4.661
NIQE↑ 5.220 6.429 5.340 5.637 5.893 5.364 4.547 5.269 5.366 5.283 6.006
UFO UIQM↑ 2.800 3.849 4.223 4.325 4.283 4.998 4.047 4.116 4.209 4.109 4.713
UICM↑ -68.97 -34.93 -20.31 -19.00 -21.88 -17.00 -33.61 -23.44 -26.55 -25.31 -2.238
NIQE↑ 4.385 4.398 4.808 4.575 4.707 5.159 4.520 4.924 5.081 5.046 6.322
UIEB UIQM↑ 2.716 3.911 4.240 4.505 4.543 4.427 4.390 4.474 4.487 4.384 4.896
UICM↑ -73.48 -32.56 -16.75 -17.43 -20.36 -19.00 -23.79 -19.78 -21.79 -21.56 5.532
NIQE↑ 6.702 8.807 6.977 7.494 6.705 5.156 4.402 6.253 4.990 5.887 6.402
@@ 4.4 Ablation Study
We conducted a detailed ablation study for two stages separately. First, we removed different components in the first stage, and fixed the second stage configuration to train the entire network. The ablation configurations are as follows: (1) -w/o DAM, representing the removal of DAM in DEN, retaining only skip connections. (2) -w/o RSB, representing the removal of feature transmission from RSB in ASN. (3) -w/o ASN, representing the removal of auxiliary supervision from ASN in the training process. Fig. 13 compares the effects of the above configurations. In the absence of DAM, the depth estimation of scene details lacks accuracy, and the removal of RSB introduces confusion in the predicted depth. Therefore, the enhanced images of two configurations suffer from excessive red channel enhancement. The depth image of -w/o ASN deviates from the underwater image, leading to inaccurate color enhancement. In Table 3, although the referenced metrics of the complete network are slightly lower than that of -w/o ASN, the advantage of the no-reference metrics still proves a significant gain of ASN.
@@
Second, we fixed the configuration of the first stage, and removed different components in the second stage to obtain the following ablation network: (1) -w/o RB, representing the removal of all RB. (2) -w/o D, representing the removal of the depth maps in DPM, retaining only the second stage. (3) -w/o Rev, representing the removal of depth map reversal. (4) -w/o RS, representing the removal of RS in the R3S transformation, directing the local branch only through the reshaped depth map. (5) -w/o DPM, representing the removal of the DPM, retaining only skip connections. In Fig. 14. (a), the color of the complete network is the most consistent with GT. When the network lacks RB or DPM, the enhanced image is still affected by color deviation. While the red channel enhancement of - w/o D, -w/o Rev and - w/o RS is insufficient, and the color distribution of -w/o RS is chaotic. In the enlarged far water region of Fig. 14. (b), only the -w/o RS and the complete network effectively counter the color deviation. In Table 4, the complete network is weaker than the -w/o RS in UIQM, but it outperforms the other configured networks in the remaining four metrics.
@@ 4.5 Extended Applications
Without adjusting any parameters, we applied the model to three challenging tasks: image segmentation [47], keypoint detection [48], and saliency detection [49, 50], while also promoting it in low-light [51, 52] and exposure [53] scenarios. Fig. 15 illustrates the results of the different tasks. The enhanced image segmentation exhibits more accurate details, and the number of keypoints increases from 49 and 27 to 73 and 56, respectively. Additionally, the saliency structure of the image is also more specific. In Fig. 16, the UVZ effectively enhances the brightness and details of the low-light image, and makes the colors of the exposed image more vivid. The above good applications further validate the effectiveness and generalization of UVZ.

### v4

@@ 3.4 Training Strategy
The UVZ training process is supervised in two stages. In the first stage, we compute losses for the outputs of DEN and ASN, respectively, which is an offline pre-training process. During the second stage, we use DEN with fixed parameters to assist DGEN in generating the enhanced image, thereby obtaining relevant losses. The first stage of training is conducted by synthesized dataset (Li et al., 2020), which contains underwater images and corresponding depth maps of five open ocean waters and coastal waters. We randomly selected 600 image pairs to train DEN and ASN for 100 epochs, and calculated the following $l_{1}$ loss:
@@
In the second stage, we selected 1200 image pairs from the CYCLE (Fabbri et al., 2018) dataset, with 1080 pairs used for 100 epochs training and the remaining 120 pairs used for subsequent testing. To ensure the enhanced images with pleasing colors and details, we formulate the total loss $L_{s}$ of DGEN using charbonnier loss and ssim loss. The former measures pixel similarity, while the latter evaluates structural similarity. The formula is as follows:
@@ 4.1 Implementation Details
To assess the performance of UVZ, we use the remaining 120 image pairs from CYCLE dataset for comparison experiments. Additionally, we used four benchmark datasets: LNRUD (Ye et al., 2022), LSUI (Peng et al., 2023), UFO (Islam et al., 2020), and UIEB (Li et al., 2019). We randomly selected 500 images from LNRUD, 400 images from LSUI, 120 images from UFO, and 80 images from UIEB for more comprehensive testing.
The proposed model was trained and tested on NVIDIA GeForce RTX 2080 Ti GPUs using the PyTorch framework. We initialized the learning rate to 0.0002 and adjusted it gradually with Adam optimizer. For the CYCLE dataset, we measure the similarity between the enhanced results and the reference images by three common referenced metrics (i.e., PSNR, SSIM and MSE). For all datasets, we evaluate the performance by three no-reference metrics (i.e., UIQM (Panetta et al., 2015), UICM and NIQE).
@@ 4.2 Comparison on reference dataset
Our UVZ was compared with some state-of-the-art methods, including four traditional methods (i.e., L2UWE (Marques and Albu, 2020), CBM (Yuan et al., 2020), TEBCF (Yuan et al., 2022), WWPE (Zhang et al., 2024c)), four deep learning methods (i.e., Water-Net (Li et al., 2019), SCNet (Fu et al., 2022b), TACL (Liu et al., 2022), Spectroformer (Khan et al., 2024)), and two depth-related methods (i.e., USUIR (Fu et al., 2022a), PUGAN (Cong et al., 2023)).
Figure 6: Enhancement comparison on the CYCLE dataset. From left to right are the raw images, the results of L2UWE (Marques and Albu, 2020), CBM (Yuan et al., 2020), TEBCF (Yuan et al., 2022), WWPE (Zhang et al., 2024c), and Water-Net (Li et al., 2019), SCNet (Fu et al., 2022b), TACL (Liu et al., 2022), USUIR (Fu et al., 2022a), PUGAN (Cong et al., 2023), Spectroformer (Khan et al., 2024), the proposed UVZ, and the ground truth. The red box represents the enlarged details.
@@
Figure 6 shows the enhancement results of all methods on the CYCLE dataset. L2UWE and USUIR improve image contrast, but noticeable color deviations persist. CBM and Water-Net have corrected colors to some extent, but the enlarged image details are blurred. TEBCF significantly enhances the details of the image but results in darkened brightness and color confusion. WWPE, TACL, PUGAN, and Spectraformer alleviate the effects of color deviation, but enlarged details are still affected by blue-green. SCNet and our UVZ effectively eliminate color deviation and enhance image visibility, with the latter highlighting more vibrant colors. Moreover, we also exhibit the histograms of the red channel in Figure 7, which is the most severely lost during underwater imaging. The curves of UVZ and GT overlap the most, indicating the closest similarity. To further observe the color distribution, we show the pixel statistics on the RGB channels in Figure 8. All methods enhance the degraded red channel to some extent, with UVZ being the most prominent. Its distribution on the red, green, and blue channels is 36.3%, 36.1%, and 29.2% respectively, which is more consistent with GT’s 35.5%, 34.5%, and 29.9%.
@@
PSNR↑ SSIM↑ MSE↓ UIQM↑ UICM↑ NIQE↑ Runtime(s)
L2UWE (Marques and Albu, 2020) 14.55 0.6440 0.06195 2.634 -76.35 5.045 3.6725
CBM (Yuan et al., 2020) 21.48 0.7461 0.01745 4.151 -31.83 6.548 0.2425
TEBCF (Yuan et al., 2022) 18.74 0.6839 0.02090 4.606 -13.71 5.137 1.6340
WWPE (Zhang et al., 2024c) 17.40 0.6412 0.02560 4.291 -16.63 6.387 0.3207
Water-Net (Li et al., 2019) 23.23 0.8196 0.01150 4.573 -16.43 5.310 0.9613
SCNet (Fu et al., 2022b) 22.48 0.8116 0.01081 4.629 -11.01 5.203 0.0520
TACL (Liu et al., 2022) 21.78 0.7701 0.01597 $4$23.29 4.325 0.0595
Spectroformer (Khan et al., 2024) 21.42 0.7910 0.01639 4.332 -20.74 5.017 0.0394
USUIR (Fu et al., 2022a) 20.55 0.7739 0.02211 4.321 -21.15 4.941 0.0255
PUGAN (Cong et al., 2023) 21.43 0.7790 0.01571 4.522 -20.74 5.253 0.0488
UVZ 26.00 0.8721 0.00435 5.044 4.781 6.586 0.0583
@@ 4.3 Comparison on no-reference dataset
We present enhancement examples from LNRUD, LSUI, UFO and UIEB datasets in Figs. 9-12. In Figure 9 and Figure 10, the color enhancement from the UVZ is the most pronounced and well-established. While other methods mitigate the effects of color deviation, blue-green color still dominates the results. We enlarged the fishes in the near-far scenes in Figure 9 (c) and Figure 10 (c), respectively. It can see that only TEBCF and UVZ effectively remove the blue-green color, but the dull color of the former reduces the attractiveness.
From Figure 11 and Figure 12, it can be observed that L2UWE exhibits bright contrast enhancement, but over-exposure affects the color. The enhanced results of TEBCF, WWPE and Water-Net have low brightness and poor clarity. CBM effectively enhances image details, while TACL, USUIR and PUGAN alleviate the impact of color deviation. However, blue-green colors still dominate in their enhanced images. Despite the significant saturation increase in SCNet and Spectroformer, they still show deviated colors in the far scenes. Benefiting from its depth perception ability, our UVZ has excellent performance in terms of color, saturation, and visibility, whether in near or far scenes.
@@
LNRUD LSUI UFO UIEB
UIQM↑ UICM↑ NIQE↑ UIQM↑ UICM↑ NIQE↑ UIQM↑ UICM↑ NIQE↑ UIQM↑ UICM↑ NIQE↑
L2UWE (Marques and Albu, 2020) 2.688 -74.80 5.322 2.882 -72.28 5.220 2.800 -68.97 4.385 2.716 -73.48 6.702
CBM (Yuan et al., 2020) 4.037 -31.29 6.684 4.215 -29.06 6.429 3.849 -34.93 4.398 3.911 -32.56 8.807
TEBCF (Yuan et al., 2022) 4.436 -16.69 5.485 4.635 -14.98 5.340 4.223 -20.31 4.808 4.240 -16.75 6.977
WWPE (Zhang et al., 2024c) 4.296 -17.15 5.550 4.439 -13.64 5.637 4.325 -19.00 4.575 4.505 -17.43 7.494
Water-Net (Li et al., 2019) 4.442 -19.72 5.694 4.607 -15.31 5.893 4.283 -21.88 4.707 4.543 -20.36 6.705
SCNet (Fu et al., 2022b) 4.452 -16.85 5.305 4.597 -11.81 5.364 4.998 -17.00 5.159 4.427 -19.00 5.156
TACL (Liu et al., 2022) 4.326 -27.65 4.465 4.460 -22.39 4.547 4.047 -33.61 4.520 4.390 -23.79 4.402
Spectroformer (Khan et al., 2024) 4.314 -20.33 5.308 4.881 -15.64 5.283 4.109 -25.31 5.046 4.384 -21.56 5.887
USUIR (Fu et al., 2022a) 4.296 -21.25 5.374 4.672 -20.31 5.269 4.116 -23.44 4.924 4.474 -19.78 6.253
PUGAN (Cong et al., 2023) 4.377 -23.03 5.133 4.542 -17.48 5.366 4.209 -26.55 5.081 4.487 -21.79 4.990
UVZ 4.619 -9.933 6.241 4.863 -4.661 6.006 4.713 -2.238 6.322 4.896 5.532 6.402
@@ 4.4 Ablation Study
We conducted a detailed ablation study for two stages separately. First, we removed different components in the first stage, and fixed the second stage configuration to train the entire network. The ablation configurations are as follows: (1) -w/o DAM, representing the removal of DAM in DEN, retaining only skip connections. (2) -w/o RSB, representing the removal of feature transmission from RSB in ASN. (3) -w/o ASN, representing the removal of auxiliary supervision from ASN in the training process. Figure 13 compares the effects of the above configurations. In the absence of DAM, the depth estimation of scene details lacks accuracy, and the removal of RSB introduces confusion in the predicted depth. Therefore, the enhanced images of two configurations suffer from excessive red channel enhancement. The depth image of -w/o ASN deviates from the underwater image, leading to inaccurate color enhancement. In Table 3, although the referenced metrics of the complete network are slightly lower than that of -w/o ASN, the advantage of the no-reference metrics still proves a significant gain of ASN.
@@
Second, we fixed the configuration of the first stage, and removed different components in the second stage to obtain the following ablation network: (1) -w/o RB, representing the removal of all RB. (2) -w/o D, representing the removal of the depth maps in DPM, retaining only the second stage. (3) -w/o Rev, representing the removal of depth map reversal. (4) -w/o RS, representing the removal of RS in the R3S transformation, directing the local branch only through the reshaped depth map. (5) -w/o DPM, representing the removal of the DPM, retaining only skip connections. In Figure 14. (a), the color of the complete network is the most consistent with GT. When the network lacks RB or DPM, the enhanced image is still affected by color deviation. While the red channel enhancement of - w/o D, -w/o Rev and - w/o RS is insufficient, and the color distribution of -w/o RS is chaotic. In the enlarged far water region of Figure 14. (b), only the -w/o RS and the complete network effectively counter the color deviation. In Table 4, the complete network is weaker than the -w/o RS in UIQM, but it outperforms the other configured networks in the remaining four metrics.
@@ 4.5 Extended Applications
Without adjusting any parameters, we applied the model to three challenging tasks: image segmentation (Lei et al., 2018), keypoint detection (Lowe, 2004), and saliency detection (Li et al., 2015; Zhang et al., 2021), while also promoting it in low-light (Lee et al., 2013; Xu et al., 2024) and exposure (Bychkovsky et al., 2011) scenarios. Figure 15 illustrates the results of the different tasks. The enhanced image segmentation exhibits more accurate details, and the number of keypoints increases from 49 and 27 to 73 and 56, respectively. Additionally, the saliency structure of the image is also more specific. In Figure 16, the UVZ effectively enhances the brightness and details of the low-light image, and makes the colors of the exposed image more vivid. The above good applications further validate the effectiveness and generalization of UVZ.
Figure 17: Comparison of pre-post detection results by the background subtraction algorithm. The top shows five consecutive frames containing notable fish movements, the middle shows the direct detection results from the algorithm, and the bottom shows the results of the algorithm after UVZ pre-processing.
To verify the gain of UVZ in real underwater task, we integrate it as a pre-processing step into the background subtraction algorithm (Sehairi et al., 2017), which is capable of capturing moving fish. Specifically, we selected the ”Dynamic Background” sequence of video frames from UNICT (Kavasidis et al., 2014; Spampinato et al., 2014) dataset as a test object, given its rich variety of marine organism movements and scene changes. As shown in Figure 17, by introducing UVZ as a pre-processing step, the algorithm’s anti-interference ability against complex background is significantly improved, resulting in fish movements appearing more prominent and clearer. The result indicates that UVZ can effectively enhance the performance of fish motion detection, making it more robust in complex underwater environments.